Given two points on the line, input the equation of the line in standard Ax + By = C. Reduce all fractional answers to lowest terms. (6,1), (5, 4)
Answer by amoresroy(361)

(y - y1) / (x - x1) = (y - y2)/(x-x2)
(y-1)/(x-6) = (y-4)/(x-5)
(y-1)(x-5) = (x-6)(y-4)
yx-5y-x+5 = xy-4x-6y+24
yx or xy cancels out
Combining like terms you get
4x-x+6y-5y = 24-5

3x +y = 19